OpenStudy (purplereign__):

HELP WITH 2 MATH QUESTIONS. FAN AND MEDAL

OpenStudy (purplereign__):

Which statement about the equation is true? 3+x=2–3x3+x=2–3x The equation has no solution. The equation has one solutions. The equation has many solutions.

OpenStudy (purplereign__):

3+X=2-3X IS FOR THE FIRST PROBLEM

OpenStudy (purplereign__):

How many solutions does the equation have?

OpenStudy (purplereign__):

0 1 many

OpenStudy (purplereign__):

4(2-5)+2=4Z-18

OpenStudy (purplereign__):

THAT'S FOR THE SECOND PROBLEM

OpenStudy (purplereign__):

for the first one i think it has one solution and for the second problem i think there is many solutions

OpenStudy (purplereign__):

@3mar

OpenStudy (awesomedude2078):

first solution only has one you're correct still working on second

OpenStudy (purplereign__):

1 solution for the second 1

OpenStudy (awesomedude2078):

yes first and second only have one

OpenStudy (awesomedude2078):

solution

OpenStudy (purplereign__):

that is wrong

OpenStudy (purplereign__):

the last one had many like i said

OpenStudy (awesomedude2078):

okay sorry can't believe everything the internet tells you right

OpenStudy (purplereign__):

well thanks for trying

OpenStudy (purplereign__):

Solve for z. 9z+10=1−3(5−3z)9z+10=1−3(5−3z) z=−4z=−4 z=−2z=−2 all real numbers no solution

OpenStudy (awesomedude2078):

you're welcome sorry i messed up

OpenStudy (purplereign__):

9z+10=1-3(5-3z)

OpenStudy (purplereign__):

that's the right problem for the question i just posted

OpenStudy (purplereign__):

I think it's all real numbers

OpenStudy (awesomedude2078):

i'd say no solution but, i don't know if it's right

OpenStudy (purplereign__):

hich statement about the equation is true? x + 2 = x + 4 The equation has no solution. The equation has one solutions. The equation has many solutions.

OpenStudy (purplereign__):

How many solutions does the equation have? 3(d +11) = 6(d + 33) 0 1 many

OpenStudy (awesomedude2078):

first one no solution because it would equal 0=2

OpenStudy (purplereign__):

ok thanks

OpenStudy (purplereign__):

now the other 2 im guessing one solution and 0

OpenStudy (retireed):

x + 2 = x + 4 2 = 4 NO Solution

OpenStudy (purplereign__):

are u 100% sure

OpenStudy (purplereign__):

and what about the last one?

OpenStudy (awesomedude2078):

one i think unless i'm wrong again

OpenStudy (retireed):

3(d +11) = 6(d + 33) 3d +33 = 6d +198 One Solution -165 = 3d -165/3 = d

OpenStudy (purplereign__):

im gonna submit it and see

OpenStudy (purplereign__):

you guys were right